# From Pixels to Titles: Video Game Identification by Screenshots using Convolutional Neural Networks

### v2

@@ 4 CNN Comparison
Regarding the systems, the best accuracy is achieved with the Xbox Series using EfficientNetB3 ($0.9714$). However, it is worth noticing that this system only had 37 screenshots from a total of five games. Its results were markedly worse with DenseNet169 ($0.6071$), for example. The Atari 2600 is the system with the best average accuracy ($0.8942$). This is likely related to the simpler graphics of this second-generation console compared to newer systems. Most games for the Atari 2600 do not exhibit significant screen variation.
@@
In simpler tasks, smaller architectures often perform as well as larger ones. It’s valuable to consider these smaller networks when selecting the optimal architecture for a task because if a smaller network can yield comparable results in less computational time, there’s no justification for employing a larger one. This rationale led to the inclusion of MobileNet and EfficientNetB0 in this comparison. However, in the context of video game detection by screenshot, it became evident that larger networks outperformed the smaller ones.
5 Alternative Initial Weights
@@
To conduct these simulations, Arcade screenshots were obtained from the Moby Games Database using the same criteria applied in sourcing screenshots from home console systems. Out of 3,125 games and 24,714 screenshots, 1,633 games and 24,235 screenshots were selected based on the criterion of ‘at least five screenshots per game.’ This Arcade dataset holds particular significance due to the inclusion of games contemporary to multiple home console generations, thereby presenting screenshots with highly diverse graphics.
@@
From the architectures that exhibited the best performance in the previous section — DenseNet169, EfficientNetB2, and EfficientNetB3 — each was trained using the entire Arcade dataset, utilizing identical parameters as outlined earlier. The weights obtained from training on the Arcade dataset were subsequently employed as initial weights for training these architectures with screenshots from each of the 22 home console systems.
@@
Tables 4, 5, and 6 display the accuracy achieved and the epochs required to train each network, using both the ImageNet and Arcade weights. For DenseNet169, employing the Arcade weights resulted in improved accuracy for only 6 out of the 22 systems. However, training times decreased for 21 of the 22 systems. Overall, while the average accuracy slightly decreased from $0.7446$ to $0.7432$, the average number of epochs needed to train the network decreased from $24.2$ to $19.6$.
Weights ImageNet Arcade
System Epochs Accuracy Epochs Accuracy
Atari 2600 22.6 0.8939 21.8 0.8971
NES 22.8 0.6719 14.4 0.7091
Master System 24.6 0.7452 16.2 0.7454
PC Engine 25.8 0.7989 17.0 0.8016
Mega Drive 22.8 0.7307 14.2 0.7252
Super Nintendo 23.8 0.7247 14.2 0.7209
Sega Saturn 23.0 0.7494 17.6 0.7211
PlayStation 23.0 0.7079 14.4 0.6838
Nintendo 64 29.8 0.8327 19.8 0.7988
Dreamcast 24.2 0.7513 19.0 0.7261
PlayStation 2 26.2 0.7273 19.8 0.7170
GameCube 25.0 0.8233 19.0 0.7942
Xbox 24.0 0.8439 20.0 0.8177
Xbox 360 24.6 0.7859 20.4 0.7625
PlayStation 3 23.2 0.7073 18.0 0.6767
Wii 23.2 0.8337 19.2 0.8102
Wii U 29.0 0.7264 22.4 0.7574
PlayStation 4 22.6 0.6151 17.6 0.5740
Xbox One 20.2 0.6458 17.0 0.6379
Nintendo Switch 25.0 0.8001 21.2 0.7869
Xbox Series 23.6 0.6071 50.0 0.8429
PlayStation 5 23.0 0.6580 17.2 0.6447
Average 24.2 0.7446 19.6 0.7432
@@
Table 4: Accuracy obtained and epochs needed to train the DenseNet169 architecture in each of the 22 screenshots datasets, with both the ImageNet weights and the Arcade dataset weights as the initial weights. Each model is executed five times following the Cross Validation approach. The highest accuracy and the lowest number of epochs for each dataset are highlighted in bold.
@@
For EfficientNetB2 (Table 5), employing the Arcade weights resulted in improved accuracy for 19 of the 22 systems. Additionally, training times decreased for 20 of the 22 systems. Overall, the average accuracy increased from $0.7343$ to $0.7552$, while the average number of epochs needed to train the network decreased from $24.4$ to $19.9$
Weights ImageNet Arcade
System Epochs Accuracy Epochs Accuracy
Atari 2600 31.6 0.8920 22.4 0.9022
NES 16.6 0.6876 13.4 0.6935
Master System 22.0 0.7418 15.6 0.7489
PC Engine 21.6 0.7834 15.0 0.7996
Mega Drive 21.0 0.7246 13.4 0.7286
Super Nintendo 18.8 0.7142 13.6 0.7153
Sega Saturn 24.6 0.6841 19.0 0.7369
PlayStation 19.6 0.7227 15.8 0.7141
Nintendo 64 30.8 0.7681 21.4 0.8205
Dreamcast 24.4 0.6735 20.0 0.7499
PlayStation 2 20.4 0.7126 18.2 0.7354
GameCube 25.0 0.7540 18.8 0.8190
Xbox 36.8 0.7929 29.2 0.8389
Xbox 360 22.0 0.7623 20.4 0.7863
PlayStation 3 19.0 0.6890 17.8 0.6901
Wii 23.2 0.7674 19.2 0.8265
Wii U 25.8 0.7006 26.8 0.7677
PlayStation 4 20.2 0.6273 15.2 0.5838
Xbox One 19.4 0.5858 16.6 0.6525
Nintendo Switch 25.0 0.7811 19.2 0.8013
Xbox Series 40.6 0.9714 42.8 0.8429
PlayStation 5 29.4 0.6188 24.6 0.6602
Average 24.4 0.7343 19.9 0.7552
@@
Table 5: Accuracy obtained and epochs needed to train the EfficientNetB2 architecture in each of the 22 screenshots datasets, with both the ImageNet weights and the Arcade dataset weights as the initial weights. Each model is executed five times following the Cross Validation approach. The highest accuracy and the lowest number of epochs for each dataset are highlighted in bold.
@@
Finally, with EfficientNetB3 (Table 6), employing the Arcade weights similarly led to improved accuracy for 19 of the 22 systems. Moreover, training times decreased for 20 of the 22 systems. Overall, the average accuracy increased from $0.7451$ to $0.7636$, while the average number of epochs required to train the network decreased from $23.7$ to $20.5$. It’s worth noting that the systems showing improved accuracy or training times are consistent between EfficientNetB2 and EfficientNetB3.
Weights ImageNet Arcade
System Epochs Accuracy Epochs Accuracy
Atari 2600 28.4 0.9036 24.8 0.9064
NES 15.2 0.6780 13.6 0.7214
Master System 22.0 0.7429 16.2 0.7562
PC Engine 19.2 0.7890 15.4 0.8108
Mega Drive 18.6 0.7307 13.8 0.7420
Super Nintendo 18.2 0.7081 15.4 0.7395
Sega Saturn 22.8 0.7050 15.2 0.7510
PlayStation 21.2 0.7402 15.4 0.7356
Nintendo 64 31.6 0.7776 23.8 0.8323
Dreamcast 23.0 0.6758 21.6 0.7643
PlayStation 2 21.6 0.7345 19.4 0.7531
GameCube 24.2 0.7498 23.8 0.8396
Xbox 29.8 0.7860 21.8 0.8511
Xbox 360 20.8 0.7768 20.2 0.7955
PlayStation 3 20.6 0.7182 18.0 0.7088
Wii 22.6 0.7840 18.4 0.8170
Wii U 22.0 0.7187 28.2 0.7652
PlayStation 4 20.4 0.6477 14.8 0.6060
Xbox One 24.4 0.6350 15.2 0.6596
Nintendo Switch 32.4 0.7886 22.4 0.8180
Xbox Series 40.0 0.9714 50.0 0.7571
PlayStation 5 23.0 0.6299 22.8 0.6676
Average 23.7 0.7451 20.5 0.7636
@@
Table 6: Accuracy obtained and epochs needed to train the EfficientNetB3 architecture in each of the 22 screenshots datasets, with both the ImageNet weights and the Arcade dataset weights as the initial weights. Each model is executed five times following the Cross Validation approach. The highest accuracy and the lowest number of epochs for each dataset are highlighted in bold.
@@
Table 7 displays the highest accuracy achieved for each system, showcasing the best combination of architecture and initial weights. EfficientNetB3 notably outperforms other architectures, yielding the best results in 19 out of 22 systems, with one tie alongside EfficientNetB2. EfficientNetB2 and DenseNet169 excel in two systems each. Concerning the initial weights, the “Arcade weights” account for the best results in 17 out of 22 systems, while the remaining five systems attained their highest accuracy with the ImageNet initial weights.
@@
Atari 2600 EfficientNetB3 Arcade 0.9064
NES EfficientNetB3 Arcade 0.7214
Master System EfficientNetB3 Arcade 0.7562
PC Engine EfficientNetB3 Arcade 0.8108
Mega Drive EfficientNetB3 Arcade 0.7420
Super Nintendo EfficientNetB3 Arcade 0.7395
Sega Saturn EfficientNetB3 Arcade 0.7510
PlayStation EfficientNetB3 ImageNet 0.7402
Nintendo 64 DenseNet169 ImageNet 0.8327
Dreamcast EfficientNetB3 Arcade 0.7643
PlayStation 2 EfficientNetB3 Arcade 0.7531
GameCube EfficientNetB3 Arcade 0.8396
Xbox EfficientNetB3 Arcade 0.8511
Xbox 360 EfficientNetB3 Arcade 0.7955
PlayStation 3 EfficientNetB3 Arcade 0.7182
Wii DenseNet169 ImageNet 0.8337
Wii U EfficientNetB2 Arcade 0.7677
PlayStation 4 EfficientNetB3 ImageNet 0.6477
Xbox One EfficientNetB3 Arcade 0.6596
Nintendo Switch EfficientNetB3 Arcade 0.8180
Xbox Series EfficientNetB2 / EfficientNetB3 (tie) ImageNet 0.9714
PlayStation 5 EfficientNetB3 Arcade 0.6676
Average 0.7767
Table 7: The highest accuracy for each system, achieved with the best combination of architecture and initial weights.
6 Conclusions

### v3

@@ V Comparison
Regarding the CNNs, EfficientNetV2S achieved the best accuracy in 16 of the 22 systems, as well as the best average accuracy across all systems ($77.44\%$). Regarding the systems, the best accuracy is achieved with the Xbox Series X/S using EfficientNetV2S ($100\%$). However, it is worth noticing that this system only had 37 screenshots from a total of five games. The Atari 2600 is the system with the best average accuracy ($84.64\%$) among all tested models. This is likely related to the simpler graphics of this second-generation console compared to newer systems. Most games for the Atari 2600 do not exhibit significant screen variation. On the other hand, newer systems with complex graphics like PlayStation 4, PlayStation 5 and XBox One are among those with the lowest average accuracy.
@@
In simpler tasks, smaller architectures often perform as well as larger ones. It’s valuable to consider these smaller networks when selecting the optimal architecture for a task because if a smaller network can yield comparable results in less computational time, there is no justification for employing a larger one. This rationale led to the inclusion of MobileNet and EfficientNetB0 in this comparison. However, in the context of video game detection by screenshot, it became evident that larger networks like DenseNet and the larger versions of EfficientNet outperformed the smaller ones. Therefore, it is not possible to use smaller networks without compromising accuracy.
Earlier architectures such as VGG and ResNet performed worse compared to newer models like DenseNet and EfficientNet. Notably, VGG struggled to learn the training subset in many instances, as indicated by its significantly lower accuracy and higher standard deviation relative to the other models.
Over recent years, transformer architectures have begun to outperform CNN models in many image classification tasks where CNNs once dominated. However, this is not the case for the video game identification by screenshots task. The three transformer architectures evaluated performed worse than all the CNN models except for VGG16. ViT-B16 achieved the highest average accuracy among the transformers, but only at $51.19\%$. SwinT was the best among transformers in nine of the 22 systems, though its overall average was lower at $44.06\%$. In many instances, the transformers struggled to learn the training subset, contributing to their higher standard deviation. Even when they did learn, their accuracy remained substandard compared to CNNs. It is known that transformers require larger training sets than CNNs to outperform them in most scenarios. This is likely the case for this task, where the number of screenshots per game (individuals per class) is relatively low.
TABLE IV: Accuracy achieved by the three different transformer models in each of the $22$ screenshots datasets. Each model is executed five times following the Cross Validation approach. The highest accuracy for each dataset is highlighted in bold. Standard deviations are shown in parentheses.
System ViT-B16 ViT-L32 SwinT Average
Atari 2600 72,49% (2,49%) 74,30% (2,88%) 2,19% (0,56%) 30,87% (16,48%)
NES 51,06% (7,03%) 58,17% (1,94%) 54,53% (1,38%) 51,06% (24,61%)
Master System 55,91% (5,25%) 58,06% (9,13%) 40,58% (25,99%) 55,91% (27,80%)
PC Engine 69,74% (1,21%) 57,05% (11,06%) 73,21% (3,26%) 69,74% (29,16%)
Mega Drive 52,43% (8,28%) 55,84% (10,09%) 35,07% (28,40%) 52,43% (27,54%)
Super Nintendo 53,12% (8,57%) 53,58% (11,08%) 33,66% (32,60%) 53,12% (27,90%)
Sega Saturn 52,07% (5,51%) 52,96% (2,73%) 50,63% (7,15%) 52,07% (23,80%)
PlayStation 42,89% (8,96%) 48,74% (8,98%) 54,70% (3,23%) 42,89% (24,92%)
Nintendo 64 57,03% (3,24%) 49,03% (18,10%) 57,97% (25,75%) 57,03% (30,82%)
Dreamcast 51,61% (2,58%) 50,02% (2,62%) 57,40% (8,27%) 51,61% (24,18%)
PlayStation 2 51,08% (3,23%) 47,83% (8,04%) 44,60% (22,13%) 51,08% (25,17%)
GameCube 57,76% (5,47%) 46,40% (20,30%) 22,41% (24,37%) 57,76% (23,79%)
Xbox 52,77% (9,66%) 52,28% (4,49%) 55,06% (26,55%) 52,77% (29,61%)
Xbox 360 42,65% (18,69%) 46,77% (18,25%) 59,32% (2,59%) 42,65% (29,12%)
PlayStation 3 45,32% (2,40%) 33,02% (10,75%) 46,53% (2,65%) 45,32% (19,07%)
Wii 60,66% (5,14%) 48,40% (20,21%) 43,57% (32,85%) 60,66% (30,03%)
Wii U 50,19% (3,35%) 49,94% (4,25%) 33,08% (21,20%) 50,19% (22,36%)
PlayStation 4 38,25% (6,17%) 28,60% (10,49%) 46,78% (1,69%) 38,25% (18,75%)
Xbox One 34,88% (10,77%) 32,30% (13,86%) 33,92% (17,75%) 34,88% (21,72%)
Nintendo Switch 50,57% (7,22%) 53,68% (5,90%) 40,23% (22,26%) 50,57% (25,86%)
Xbox Series X/S 52,50% (22,52%) 48,93% (7,79%) 58,21% (22,97%) 52,50% (32,08%)
PlayStation 5 31,17% (8,47%) 27,09% (1,99%) 25,57% (11,73%) 31,17% (14,97%)
Average 51,19% (7,10%) 48,77% (9,32%) 44,06% (15,70%) 49,30% (24,99%)
VI Alternative Initial Weights
@@
To conduct these experiments, we used the platform listed on MobyGames as ‘Arcade’. They have included all the arcade games in this single platform. Therefore, it contains games contemporary to multiple home console generations, with graphics of increasing complexity being released over the years. The Arcade screenshots were obtained using the same criteria applied in sourcing screenshots from home console systems. Out of 3,125 games and 24,714 screenshots, 1,633 games and 24,235 screenshots were selected based on the criterion of ‘at least five screenshots per game.’ It is worth noticing that some arcade games were later ported to home consoles, usually with simpler graphics due to hardware limitations. Despite that, convolutional layers trained in similar graphics could be more easily fine-tuned to detect the home version counterparts.
@@
The three architectures that demonstrated the best accuracies in the previous section — DenseNet201, EfficientNetB3, and EfficientNetV2S — were chosen for application in these new experiments. Each of them was trained using the entire Arcade dataset, utilizing identical parameters as outlined earlier. The weights obtained from training on the Arcade dataset were subsequently employed as initial weights for training these architectures with screenshots from each of the 22 home console systems. It is worth noting that the golden age of arcades spanned from the late 1970s to the early 1980s, and the number of new releases declined significantly afterward. In the new millennium, there are even fewer arcade game releases. Consequently, it is expected that earlier home consoles could derive greater benefit from using the Arcade weights.
@@
Tables V, VI, and VII display the accuracy achieved and the epochs required to train each network, using random weights initialization and both the ImageNet and Arcade weights. In all scenarios, random weights resulted in lower accuracy and required more epochs for convergence compared to the other weight initialization methods. Conversely, Arcade weights outperformed ImageNet weights in most scenarios, though not all. For DenseNet201 (Table V), employing the Arcade weights resulted in improved accuracy for only 10 out of the 22 systems. However, training times decreased for 20 of the 22 systems. Overall, while the average accuracy remained the same ($75.77\%$), the average number of epochs needed to train the network decreased from $24.6$ to $22.0$.
@@
TABLE V: Accuracy and epochs required to train the DenseNet201 architecture across 22 screenshot datasets, using random weight initialization, ImageNet weights, and Arcade dataset weights as initial weights. Each model is executed five times using the Cross Validation approach. The highest accuracy and the lowest number of epochs for each dataset are highlighted in bold. Standard deviations are shown in parentheses.
Weights Random ImageNet Arcade
System Epochs Accuracy Epochs Accuracy Epochs Accuracy
Atari 2600 44.6 (3.2) 80.82% (2.17%) 23.6 (1.5) 89.06% (0.53%) 17.6 (0.8) 89.99% (0.98%)
NES 25.4 (1.2) 65.46% (0.85%) 23.0 (3.9) 70.04% (1.91%) 20.8 (2.5) 70.77% (3.14%)
Master System 38.2 (3.0) 67.11% (1.32%) 24.0 (1.5) 73.74% (1.14%) 20.0 (0.9) 74.14% (1.00%)
PC Engine 38.2 (5.1) 74.90% (0.68%) 23.6 (2.1) 80.19% (0.70%) 18.4 (1.4) 80.52% (1.59%)
Mega Drive 29.2 (2.9) 67.40% (1.74%) 24.0 (3.0) 73.44% (2.53%) 21.4 (4.1) 71.48% (4.19%)
Super Nintendo 26.8 (2.1) 66.48% (1.03%) 25.8 (0.7) 74.51% (0.80%) 22.0 (2.9) 73.32% (2.36%)
Sega Saturn 37.8 (3.9) 65.17% (2.42%) 22.4 (1.5) 75.10% (1.35%) 19.2 (1.3) 83.14% (1.15%)
PlayStation 25.4 (1.7) 63.52% (2.36%) 22.6 (3.0) 71.19% (2.78%) 23.0 (3.8) 71.59% (2.44%)
Nintendo 64 46.0 (3.6) 73.93% (3.35%) 26.2 (3.0) 83.67% (1.17%) 19.2 (0.7) 83.37% (2.11%)
Dreamcast 48.8 (1.2) 63.35% (2.18%) 25.0 (2.8) 75.36% (1.90%) 20.4 (3.0) 76.06% (1.40%)
PlayStation 2 37.8 (5.5) 62.00% (2.22%) 25.2 (3.2) 74.11% (0.39%) 22.0 (2.0) 73.13% (0.47%)
GameCube 50.0 (0.0) 71.71% (1.22%) 23.6 (1.5) 82.88% (1.14%) 22.8 (2.3) 82.52% (1.13%)
Xbox 46.2 (4.0) 74.25% (1.05%) 24.0 (1.4) 84.03% (1.38%) 22.4 (1.4) 84.06% (1.62%)
Xbox 360 41.8 (5.2) 68.01% (1.03%) 23.6 (2.4) 79.43% (0.72%) 21.6 (1.5) 78.71% (0.54%)
PlayStation 3 36.8 (8.5) 52.08% (6.77%) 23.8 (2.3) 70.93% (0.90%) 21.0 (1.4) 70.96% (0.54%)
Wii 46.6 (3.6) 73.96% (1.60%) 23.4 (2.1) 83.05% (0.79%) 20.6 (1.4) 82.49% (0.71%)
Wii U 50.0 (0.0) 56.39% (2.02%) 25.6 (6.4) 73.03% (2.66%) 20.6 (2.7) 72.90% (3.46%)
PlayStation 4 26.6 (4.6) 49.18% (5.30%) 24.0 (1.9) 62.19% (0.70%) 22.4 (1.5) 61.63% (0.55%)
Xbox One 43.6 (1.9) 54.28% (1.27%) 21.4 (2.6) 64.96% (1.33%) 20.2 (2.4) 66.25% (1.41%)
Nintendo Switch 49.2 (1.6) 67.74% (2.49%) 24.8 (4.3) 79.90% (1.55%) 24.2 (4.3) 78.98% (0.61%)
Xbox Series X/S 20.6 (15.0) 32.14% (12.78%) 36.2 (16.9) 78.57% (5.98%) 42.4 (15.2) 73.57% (13.80%)
PlayStation 5 42.4 (2.7) 47.82% (1.97%) 24.6 (4.0) 67.58% (2.22%) 21.8 (2.6) 67.36% (1.29%)
Average 38.7 (3.7) 63.53% (2.63%) 24.6 (3.3) 75.77% (1.57%) 22.0 (2.7) 75.77% (2.11%)
@@
For EfficientNetB3 (Table VI), employing the Arcade weights led to improved accuracy for 19 of the 22 systems. Moreover, training times decreased for 20 of the 22 systems. Overall, the average accuracy increased from $74.51\%$ to $76.36\%$, while the average number of epochs required to train the network decreased from $23.7$ to $20.5$. This architecture demonstrated the most significant improvement using the Arcade weights.
@@
TABLE VI: Accuracy and epochs required to train the EfficientNetB3 architecture across 22 screenshot datasets, using random weight initialization, ImageNet weights, and Arcade dataset weights as initial weights. Each model is executed five times using the Cross Validation approach. The highest accuracy and the lowest number of epochs for each dataset are highlighted in bold. Standard deviations are shown in parentheses.
Weights Random ImageNet Arcade
System Epochs Accuracy Epochs Accuracy Epochs Accuracy
Atari 2600 47.4 (3.6) 77.47% (2.28%) 28.4 (6.1) 90.36% (1.52%) 24.8 (6.7) 90.64% (1.25%)
NES 22.0 (0.9) 53.75% (2.20%) 15.2 (1.5) 67.80% (4.17%) 13.6 (0.5) 72.14% (0.74%)
Master System 22.0 (1.1) 48.61% (3.19%) 22.0 (1.7) 74.29% (1.23%) 16.2 (2.7) 75.62% (0.97%)
PC Engine 29.0 (3.2) 64.42% (1.95%) 19.2 (1.0) 78.90% (0.84%) 15.4 (0.5) 81.08% (0.71%)
Mega Drive 21.2 (0.4) 54.28% (2.34%) 18.6 (2.9) 73.07% (4.22%) 13.8 (0.8) 74.20% (0.60%)
Super Nintendo 21.8 (1.2) 52.30% (1.14%) 18.2 (3.7) 70.81% (6.09%) 15.4 (2.6) 73.95% (1.53%)
Sega Saturn 22.6 (3.2) 45.93% (5.07%) 22.8 (2.8) 70.50% (0.98%) 15.2 (0.8) 75.10% (1.60%)
PlayStation 22.2 (0.4) 53.81% (1.52%) 21.2 (3.1) 74.02% (1.03%) 15.4 (1.7) 73.56% (0.63%)
Nintendo 64 34.0 (10.5) 58.18% (1.26%) 31.6 (10.3) 77.76% (0.90%) 23.8 (5.6) 83.23% (1.66%)
Dreamcast 35.0 (6.8) 48.44% (2.78%) 23.0 (1.7) 67.58% (0.97%) 21.6 (2.2) 76.43% (1.52%)
PlayStation 2 23.4 (0.8) 52.47% (1.46%) 21.6 (3.3) 73.45% (2.52%) 19.4 (2.0) 75.31% (1.21%)
GameCube 30.2 (4.0) 58.80% (1.22%) 24.2 (2.9) 74.98% (1.40%) 23.8 (6.0) 83.96% (0.96%)
Xbox 27.6 (2.6) 63.89% (1.43%) 29.8 (5.7) 78.60% (2.33%) 21.8 (3.7) 85.11% (1.50%)
Xbox 360 25.6 (3.8) 52.29% (3.92%) 20.8 (0.8) 77.68% (0.86%) 20.2 (1.6) 79.55% (0.65%)
PlayStation 3 24.0 (0.9) 45.97% (1.19%) 20.6 (3.1) 71.82% (0.92%) 18.0 (2.1) 70.88% (0.47%)
Wii 35.6 (6.7) 62.13% (1.53%) 22.6 (1.4) 78.40% (1.64%) 18.4 (4.4) 81.70% (1.11%)
Wii U 33.2 (17.6) 27.48% (16.66%) 22.0 (4.2) 71.87% (2.93%) 28.2 (4.8) 76.52% (4.23%)
PlayStation 4 22.2 (1.7) 40.53% (0.99%) 20.4 (1.6) 64.77% (1.53%) 14.8 (1.6) 60.60% (1.69%)
Xbox One 22.4 (1.0) 40.76% (1.17%) 24.4 (1.4) 63.50% (2.92%) 15.2 (1.5) 65.96% (1.38%)
Nintendo Switch 27.4 (2.7) 44.24% (1.35%) 32.4 (7.2) 78.86% (1.34%) 22.4 (3.4) 81.80% (1.09%)
Xbox Series X/S 47.6 (4.8) 16.07% (4.52%) 40.0 (13.1) 97.14% (5.71%) 50.0 (0.0) 75.71% (16.66%)
PlayStation 5 24.6 (1.5) 35.75% (3.38%) 23.0 (1.1) 62.99% (2.57%) 22.8 (5.8) 66.76% (1.68%)
Average 28.2 (3.6) 49.89% (2.84%) 23.7 (3.7) 74.51% (2.21%) 20.5 (2.8) 76.36% (1.99%)
@@
Finally, with EfficientNetV2S (Table VII), using the Arcade weights led to improved accuracy in only 9 out of the 22 systems. However the average accuracy still increased from $77.44\%$ to $77.63\%$. Training times decreased for 17 of the 22 systems, and the average number of epochs required to train the network decreased from $26.9$ to $24.5$.
@@
TABLE VII: Accuracy and epochs required to train the EfficientNetV2S architecture across 22 screenshot datasets, using random weight initialization, ImageNet weights, and Arcade dataset weights as initial weights. Each model is executed five times using the Cross Validation approach. The highest accuracy and the lowest number of epochs for each dataset are highlighted in bold. Standard deviations are shown in parentheses.
Weights Random ImageNet Arcade
System Epochs Accuracy Epochs Accuracy Epochs Accuracy
Atari 2600 42.0 (6.2) 82.31% (1.77%) 31.8 (3.7) 88.92% (2.74%) 28.8 (5.8) 88.78% (0.82%)
NES 23.0 (0.6) 60.43% (2.11%) 20.6 (1.9) 72.51% (2.40%) 17.4 (0.5) 69.33% (2.64%)
Master System 27.8 (3.0) 61.61% (1.32%) 26.0 (1.4) 76.20% (1.19%) 24.2 (1.2) 74.45% (0.81%)
PC Engine 25.6 (1.5) 68.58% (2.00%) 25.6 (1.6) 80.52% (0.94%) 23.6 (2.6) 79.33% (0.46%)
Mega Drive 23.6 (1.4) 59.26% (1.90%) 27.2 (1.3) 74.56% (3.22%) 20.0 (2.1) 73.11% (1.87%)
Super Nintendo 22.0 (0.6) 59.10% (1.38%) 21.6 (2.7) 72.89% (1.72%) 17.6 (2.7) 71.88% (2.25%)
Sega Saturn 27.0 (3.0) 56.83% (1.96%) 31.0 (3.6) 75.91% (1.11%) 31.8 (8.2) 83.14% (1.26%)
PlayStation 23.2 (1.6) 59.38% (1.88%) 22.0 (3.0) 74.77% (2.49%) 19.4 (2.2) 73.62% (2.20%)
Nintendo 64 31.0 (1.5) 67.11% (2.21%) 33.2 (5.3) 82.28% (3.19%) 26.4 (2.9) 85.23% (1.89%)
Dreamcast 32.6 (6.8) 59.80% (1.18%) 41.2 (6.4) 74.99% (1.83%) 25.6 (3.7) 78.86% (1.45%)
PlayStation 2 24.2 (1.2) 55.53% (1.81%) 25.4 (1.4) 75.14% (3.29%) 25.4 (2.7) 76.73% (0.63%)
GameCube 30.0 (2.0) 63.74% (2.13%) 28.0 (2.4) 81.48% (2.30%) 32.2 (7.6) 85.00% (0.75%)
Xbox 32.2 (4.7) 69.38% (1.94%) 35.0 (5.8) 84.45% (1.32%) 25.8 (4.4) 86.42% (1.45%)
Xbox 360 27.4 (2.6) 60.78% (1.36%) 25.8 (2.3) 80.25% (1.17%) 24.0 (1.4) 79.01% (0.44%)
PlayStation 3 27.0 (1.8) 51.48% (1.03%) 22.8 (2.8) 71.98% (3.04%) 23.4 (2.2) 72.47% (1.08%)
Wii 29.2 (2.9) 69.00% (1.75%) 29.4 (8.5) 83.96% (0.69%) 24.6 (3.0) 83.41% (0.53%)
Wii U 26.8 (3.2) 51.87% (1.81%) 24.8 (3.3) 74.97% (2.43%) 20.4 (5.1) 74.58% (3.80%)
PlayStation 4 23.8 (1.6) 46.96% (2.05%) 21.2 (2.4) 63.17% (2.88%) 20.0 (1.8) 62.53% (1.27%)
Xbox One 27.2 (2.7) 47.73% (2.49%) 25.0 (2.4) 66.25% (2.17%) 22.2 (1.9) 66.04% (1.09%)
Nintendo Switch 23.0 (1.7) 53.22% (1.65%) 34.2 (8.8) 80.47% (1.70%) 33.2 (9.0) 80.59% (1.50%)
Xbox Series X/S 29.8 (16.7) 32.86% (21.39%) 15.6 (1.7) 100.00% (0.00%) 27.8 (6.5) 94.64% (6.59%)
PlayStation 5 22.2 (1.7) 45.74% (2.83%) 23.4 (2.9) 68.10% (2.52%) 25.4 (5.2) 68.62% (1.74%)
Average 27.3 (3.1) 58.30% (2.73%) 26.9 (3.4) 77.44% (2.02%) 24.5 (3.8) 77.63% (1.66%)
@@
Table VIII displays the highest accuracy achieved for each system, showcasing the best combination of architecture and initial weights. EfficientNetV2S notably outperforms other architectures, yielding the best results in 15 out of 22 systems, with one tie alongside DenseNet201. EfficientNetB3 excels in six systems, while DenseNet201 is the best in two system, one of them alongside EfficientNetV2S. Concerning the initial weights, the “Arcade weights” account for the best results in 12 out of 22 systems, while the other ten systems attained their highest accuracy with the ImageNet initial weights. Contrary to expectations, systems benefiting from the use of arcade weights span multiple generations. This indicates that the complexity of arcade graphics is adequate for producing better initial weights for various generations of home systems.
TABLE VIII: The highest accuracy for each system, achieved with the best combination of architecture and initial weights.
@@
Atari 2600 EfficientNetB3 Arcade 90.64%
NES EfficientNetV2S ImageNet 72.51%
Master System EfficientNetV2S ImageNet 76.20%
PC Engine EfficientNetB3 Arcade 81.08%
Mega Drive EfficientNetV2S ImageNet 74.56%
Super Nintendo DenseNet201 ImageNet 74.51%
Sega Saturn DenseNet201 EfficientNetV2S Arcade 83.14%
PlayStation EfficientNetV2S ImageNet 74.77%
Nintendo 64 EfficientNetV2S Arcade 85.28%
Dreamcast EfficientNetV2S Arcade 78.86%
PlayStation 2 EfficientNetV2S Arcade 76.73%
GameCube EfficientNetB3 Arcade 83.96%
Xbox EfficientNetV2S Arcade 86.42%
Xbox 360 EfficientNetV2S ImageNet 80.25%
PlayStation 3 EfficientNetV2S Arcade 72.47%
Wii EfficientNetV2S ImageNet 83.96%
Wii U EfficientNetB3 Arcade 76.52%
PlayStation 4 EfficientNetB3 ImageNet 64.77%
Xbox One EfficientNetV2S ImageNet 66.25%
Nintendo Switch EfficientNetB3 Arcade 81.80%
Xbox Series EfficientNetV2S ImageNet 100.00%
PlayStation 5 EfficientNetV2S Arcade 68.62%
Average 78.79%
VII Limitations
One of the limitations of the current approach is that there is a separate network for each console system, which means the console system must be known prior to title identification. In future works, this can be addressed in different ways. The most straightforward approach would be to treat each game on each system as a single entity. The drawback is that the resulting network would have more than a hundred thousand outputs. Another approach would be to build a network to detect the console system first, and then another network specialized for that system, similar to those in this paper, would identify the title. The drawback in this scenario is that identifying the system could be challenging, especially with newer systems that do not have distinctive color palettes or graphics complexity limitations like older systems. Finally, a third approach would be to consider each game as an entity regardless of the system, considering that some games are available on multiple systems. A network with two parallel output layers could be used to identify the title and the system simultaneously. Given a screenshot, the probabilities of each system and each title could be used together to identify the title and system among the existing combinations, i.e., combinations of titles and systems that do not exist would be ruled out, and the existing pair with the highest combined probabilities would be chosen. However, the size of the output layer would be almost as large as the first approach.
Another limitation of the presented approach is that adding new game titles requires adding an additional output node to the output layer. Once a node is added, the new connection weights must be learned, so the network would need at least a few more epochs of training. On the other hand, the convolutional weights, especially in the first layers, may not change significantly with the addition of new titles. Therefore, their existing weights could be fine-tuned while the new connections are trained.
VIII Conclusions